# Towards Anomaly-Aware Pre-Training and Fine-Tuning for Graph Anomaly Detection

### v2

@@ G.2.1 Standard GNN Architectures
  • GCN [34] employs a convolutional operation on the graph to propagate information from each node to its neighboring nodes, enabling the network to learn a representation for each node based on its local neighborhood.
@@
  • GIN [69] is designed to capture the structural properties of a graph while preserving graph isomorphism. Specifically, it generates identical embeddings for graphs that are structurally identical, regardless of permutations in node labels.
@@
  • GAT [62] incorporates an attention mechanism, assigning different levels of importance to nodes during the information aggregation process. This allows the model to focus on the most relevant nodes within a neighborhood.
@@
  • ACM [45] leverages low-, high, and full-pass spectral filters and an attention-based mixing mechanism to adaptively extract richer localized information for diverse node heterophily situations.
@@
  • FAGCN [4] adaptively integrates low-frequency and high-frequency signals through a self-gating mechanism. This approach enhances the model’s ability to handle both homophilic and heterophilic networks.
@@
  • AdaGNN [18] leverages an adaptive frequency response filter to capture the varying importance of different frequency components for node representation learning. This approach improves the expressiveness of the model and alleviates the over-smoothing problem.
@@
  • BernNet [25] provides a robust framework for designing and learning arbitrary graph spectral filters. It uses an order-K Bernstein polynomial approximation to estimate filters over the normalized Laplacian spectrum of a graph.
G.2.2 GNNs Specialized for GAD
@@
  • GAS [37] is a highly scalable method for detecting spam reviews. It extends GCN to handle heterogeneous and heterophilic graphs and adapts to the graph structure of specific GAD applications using the KNN algorithm.
@@
  • DCI [67] reduces inconsistencies between node behavior patterns and label semantics, and captures intrinsic graph properties within concentrated feature spaces by clustering the graph into multiple segments.
@@
  • PCGNN [40] uses a label-balanced sampler to select nodes and edges for training, ensuring a balanced label distribution in the induced subgraph. Additionally, it employs a learnable, parameterized distance function to select neighbors, filtering out redundant links while adding beneficial ones for improved fraud prediction.
@@
  • AMNet [6] captures both low- and high-frequency signals by stacking multiple BernNets, adaptively combining signals from different frequency ranges.
@@
  • BWGNN [60] addresses the "right-shift" phenomenon in graph anomalies, where spectral energy distribution shifts from low to high frequencies. It uses a Beta kernel to address high-frequency anomalies through flexible, spatially- and spectrally-localized band-pass filters.
@@
  • GHRN [21] targets the heterophily problem in the spectral domain for graph anomaly detection. This method prunes inter-class edges to highlight and delineate the graph’s high-frequency components.
@@
  • ConsisGAD [8] focuses on graph anomaly detection with limited supervision. It incorporates learnable data augmentation to utilize the abundance of unlabeled data for consistency training.
@@
  • SpaceGNN [17] integrates learnable space projection, distance-aware propagation, and multiple space ensemble modules to leverage the benefits of different spaces (Euclidean, hyperbolic, and spherical) for node anomaly detection with extremely limited labels.
@@
  • XGBGraph [59] first aggregates features from neighboring nodes to enhance the representation of each node, and then uses XGBoost [9] to classify nodes as normal or anomalous. This approach leverages the robustness and efficiency of tree ensembles while incorporating graph structure to improve anomaly detection performance.
G.2.3 Graph Pre-Training Methods
@@
  • DGI [63] learns representations by maximizing the mutual information between node representations and a global summary representation.
@@
  • GRACE [63] learns node representations by pulling together the representations of the same node (positive pairs) across two augmented views, while pushing apart the representations of different nodes (negative pairs) across both views.
@@
  • GraphMAE [29] is a masked graph auto-encoder that focuses on feature reconstruction using both a masking strategy and scaled cosine error.
@@
  • BGRL [61] employs asymmetric architectures to learn node representations by predicting alternative augmentations of the input graph and maximizing the similarity between these predictions and their corresponding targets.
@@
  • G-BT [3] utilizes a cross-correlation-based loss function to reduce redundancy in the learned representations, which enjoys fewer hyperparameters and significantly reduced computation time.
@@
  • SSGE [43] minimizes the distance between the distribution of learned representations and an isotropic Gaussian distribution, promoting the uniformity of node representations.
@@
  • PolyGCL [7] addresses heterophilic challenges in graph pre-training by using polynomial filters as encoders and incorporating a combined linear objective between low- and high-frequency components in the spectral domain.
@@
  • BWDGI pre-trains the state-of-the-art GAD backbone BWGNN [60] using DGI [63] as the pretext objective.
G.3 Evaluation Protocols
@@
Following GADBench [59], we evaluate performance using three popular metrics: the Area Under the Receiver Operating Characteristic Curve (AUROC), the Area Under the Precision-Recall Curve (AUPRC) calculated via average precision, and the Recall score within the top-$K$ predictions (Rec@K). Here, $K$ corresponds to the number of anomalies in the test set. For all metrics, anomalies are treated as the positive class, with higher scores indicating better model performance. To closely simulate real-world scenarios with limited supervision, we standardize the training/validation set across all datasets to include 100 labels — 20 positive (abnormal) and 80 negative (normal) labels [59]. To ensure the robustness of our findings, we perform 10 random splits, as provided by GADBench, on each dataset and report the average performance.
G.4 Implementation Details
@@
We use the implementations of all baseline methods provided by GADBench [59] or the respective authors. Our APF model is implemented using PyTorch and the Deep Graph Library (DGL) [65]. Experiments are conducted on a Linux server equipped with an Intel(R) Xeon(R) Gold 6248 CPU @ 2.50GHz and a 32GB NVIDIA Tesla V100 GPU.
@@
During the pre-training phase, each model is trained for up to 800 epochs using the Adam optimizer [33], with a patience of 20. Hyperparameters are tuned as follows: filter order $\in\{2,3\}$, learning rate $\in\{0.01,0.001,0.0001\}$, representation dimension $\in\{32,64\}$, activation function $\in\{\text{ReLU},\text{ELU},\text{PReLU},\text{Tanh}\}$, and normalization $\in\{\text{none},\text{batch},\text{layer}\}$. For efficiency, we extract 1-hop sub-trees instead of 2-hop ones on denser or larger datasets Amazon, T-Finance, and T-Social.
@@
During the fine-tuning phase, a 2-layer MLP classifier is trained for up to 500 epochs using the Adam optimizer [33], with a learning rate of 0.01 and weight decay selected from $\{0.0,0.01,0.0001\}$. The classifier with the highest validation AUROC score is selected for testing. The hyperparameters $p_{n}$ and $p_{a}$ are varied within the range of 0.0 to 1.0.
G.5 Additional Experiments
Efficiency Comparison.
@@ G.5 Additional Experiments
Under Varying Supervision.
This section evaluates the performance of APF under varying levels of supervision by modifying the number of labeled abnormal nodes. Following [59], the number of labeled normal nodes is set to four times the number of labeled abnormal nodes. We present the results in terms of AUPRC, AUROC, and Rec@K in Figures 7, 8, and 9, respectively. As expected, performance generally improves across all methods as the number of labeled nodes increases. Notably, APF delivers consistent improvements over baseline pre-training methods and surpasses the state-of-the-art GAD models, even with only 5 labeled abnormal nodes. This highlights the effectiveness of our approach in addressing GAD with limited supervision.
@@
Additional Results for Model Comparison.
@@
Table 5: Comparison of AUROC for each model. "-" denotes "out of memory". The best and runner-up models are bolded and underlined.
@@
Table 6: Comparison of Rec@K for each model. "-" denotes "out of memory". The best and runner-up models are bolded and underlined.
@@
Additional Results for Homophily Disparity.
We provide additional results in terms of AUROC and Rec@K for homophily disparity in Figure 10 and Figure 11, respectively.
@@
Additional Results for Hyperparameter Analysis.
We provide additional results in terms of AUROC and Rec@K for hyperparameter analysis in Figure 12 and Figure 13, respectively.
(a) Weibo
(b) Amazon
(c) YelpChi
(d) T-Finance
(a) Weibo
(b) Amazon
(c) YelpChi
(d) T-Finance
NeurIPS Paper Checklist

### v3

@@ H.2.1 Standard GNN Architectures
  • GCN Kipf and Welling (2017) employs a convolutional operation on the graph to propagate information from each node to its neighboring nodes, enabling the network to learn a representation for each node based on its local neighborhood.
@@
  • GIN Xu et al. (2019) is designed to capture the structural properties of a graph while preserving graph isomorphism. Specifically, it generates identical embeddings for graphs that are structurally identical, regardless of permutations in node labels.
@@
  • GAT Veličković et al. (2018) incorporates an attention mechanism, assigning different levels of importance to nodes during the information aggregation process. This allows the model to focus on the most relevant nodes within a neighborhood.
@@
  • ACM Luan et al. (2022) leverages low-, high, and full-pass spectral filters and an attention-based mixing mechanism to adaptively extract richer localized information for diverse node heterophily situations.
@@
  • FAGCN Bo et al. (2021) adaptively integrates low-frequency and high-frequency signals through a self-gating mechanism. This approach enhances the model’s ability to handle both homophilic and heterophilic networks.
@@
  • AdaGNN Dong et al. (2021) leverages an adaptive frequency response filter to capture the varying importance of different frequency components for node representation learning. This approach improves the expressiveness of the model and alleviates the over-smoothing problem.
@@
  • BernNet He et al. (2021) provides a robust framework for designing and learning arbitrary graph spectral filters. It uses an order-K Bernstein polynomial approximation to estimate filters over the normalized Laplacian spectrum of a graph.
H.2.2 GNNs Specialized for GAD
@@
  • GAS Li et al. (2019) is a highly scalable method for detecting spam reviews. It extends GCN to handle heterogeneous and heterophilic graphs and adapts to the graph structure of specific GAD applications using the KNN algorithm.
@@
  • DCI Wang et al. (2021) reduces inconsistencies between node behavior patterns and label semantics, and captures intrinsic graph properties within concentrated feature spaces by clustering the graph into multiple segments.
@@
  • PCGNN Liu et al. (2021a) uses a label-balanced sampler to select nodes and edges for training, ensuring a balanced label distribution in the induced subgraph. Additionally, it employs a learnable, parameterized distance function to select neighbors, filtering out redundant links while adding beneficial ones for improved fraud prediction.
@@
  • AMNet Chai et al. (2022) captures both low- and high-frequency signals by stacking multiple BernNets, adaptively combining signals from different frequency ranges.
@@
  • BWGNN Tang et al. (2022) addresses the ”right-shift” phenomenon in graph anomalies, where spectral energy distribution shifts from low to high frequencies. It uses a Beta kernel to address high-frequency anomalies through flexible, spatially- and spectrally-localized band-pass filters.
@@
  • GHRN Gao et al. (2023) targets the heterophily problem in the spectral domain for graph anomaly detection. This method prunes inter-class edges to highlight and delineate the graph’s high-frequency components.
@@
  • ConsisGAD Chen et al. (2024b) focuses on graph anomaly detection with limited supervision. It incorporates learnable data augmentation to utilize the abundance of unlabeled data for consistency training.
@@
  • SpaceGNN Dong et al. (2025) integrates learnable space projection, distance-aware propagation, and multiple space ensemble modules to leverage the benefits of different spaces (Euclidean, hyperbolic, and spherical) for node anomaly detection with extremely limited labels.
@@
  • XGBGraph Tang et al. (2023) first aggregates features from neighboring nodes to enhance the representation of each node, and then uses XGBoost Chen and Guestrin (2016) to classify nodes as normal or anomalous. This approach leverages the robustness and efficiency of tree ensembles while incorporating graph structure to improve anomaly detection performance.
H.2.3 Graph Pre-Training Methods
@@
  • DGI Veličković et al. (2019) learns representations by maximizing the mutual information between node representations and a global summary representation.
@@
  • GRACE Veličković et al. (2019) learns node representations by pulling together the representations of the same node (positive pairs) across two augmented views, while pushing apart the representations of different nodes (negative pairs) across both views.
@@
  • GraphMAE Hou et al. (2022) is a masked graph auto-encoder that focuses on feature reconstruction using both a masking strategy and scaled cosine error.
@@
  • BGRL Thakoor et al. (2022) employs asymmetric architectures to learn node representations by predicting alternative augmentations of the input graph and maximizing the similarity between these predictions and their corresponding targets.
@@
  • G-BT Bielak et al. (2022) utilizes a cross-correlation-based loss function to reduce redundancy in the learned representations, which enjoys fewer hyperparameters and significantly reduced computation time.
@@
  • SSGE Liu et al. (2024c) minimizes the distance between the distribution of learned representations and an isotropic Gaussian distribution, promoting the uniformity of node representations.
@@
  • PolyGCL Chen et al. (2024a) addresses heterophilic challenges in graph pre-training by using polynomial filters as encoders and incorporating a combined linear objective between low- and high-frequency components in the spectral domain.
@@
  • BWDGI pre-trains the state-of-the-art GAD backbone BWGNN Tang et al. (2022) using DGI Veličković et al. (2019) as the pretext objective.
H.3 Evaluation Protocols
@@
Following GADBench Tang et al. (2023), we evaluate performance using three popular metrics: the Area Under the Receiver Operating Characteristic Curve (AUROC), the Area Under the Precision-Recall Curve (AUPRC) calculated via average precision, and the Recall score within the top-$K$ predictions (Rec@K). Here, $K$ corresponds to the number of anomalies in the test set. We prioritize these threshold-independent (AUROC, AUPRC) and rank-based (Rec@K) metrics over the F1 score to avoid the instability associated with selecting decision thresholds under label-scarce conditions. For all metrics, anomalies are treated as the positive class, with higher scores indicating better model performance. To closely simulate real-world scenarios with limited supervision, we strictly adhere to the semi-supervised setting defined in GADBench. Accordingly, we standardize the training/validation set across all datasets to include 100 labels — 20 positive (abnormal) and 80 negative (normal) labels Tang et al. (2023). This specific configuration ensures our results are directly comparable to the semi-supervised benchmarks reported in the GADBench Appendix (Table 11). To ensure the robustness of our findings, we perform 10 random splits, as provided by GADBench, on each dataset and report the average performance.
H.4 Implementation Details
@@
We use the implementations of all baseline methods provided by GADBench Tang et al. (2023) or the respective authors. Our APF model is implemented using PyTorch and the Deep Graph Library (DGL) Wang et al. (2019). Experiments are conducted on a Linux server equipped with an Intel(R) Xeon(R) Gold 6248 CPU @ 2.50GHz and a 32GB NVIDIA Tesla V100 GPU.
@@
During the pre-training phase, each model is trained for up to 800 epochs using the Adam optimizer Kingma and Ba (2015), with a patience of 20. Hyperparameters are tuned as follows: filter order $\in\{2,3\}$, learning rate $\in\{0.01,0.001,0.0001\}$, representation dimension $\in\{32,64\}$, activation function $\in\{\text{ReLU},\text{ELU},\text{PReLU},\text{Tanh}\}$, and normalization $\in\{\text{none},\text{batch},\text{layer}\}$. For efficiency, we extract 1-hop subgraphs instead of 2-hop ones on denser or larger datasets Amazon, T-Finance, and T-Social.
@@
During the fine-tuning phase, a 2-layer MLP classifier is trained for up to 500 epochs using the Adam optimizer Kingma and Ba (2015), with a learning rate of 0.01 and weight decay selected from $\{0.0,0.01,0.0001\}$. The classifier with the highest validation AUROC score is selected for testing. Regarding the regularization hyperparameters, while we conduct a full grid search ($p_{n},p_{a}\in[0,1]$) for the sensitivity analysis, we adopt an efficient strategy for practical tuning: we fix $p_{n}$ to a high default value (e.g., $0.9$ or $1.0$) and perform a small search for $p_{a}$ (e.g., $\{0.0,0.1,0.2,0.3,0.4\}$), subject to $p_{a}\leq p_{n}$. Our implementation codes are available at https://github.com/Cloudy1225/APF.
Appendix I Additional Experiments
I.1 Pretraining-only for Unsupervised GAD
To further verify the effectiveness of our pre-training, we also investigate its performance under a purely unsupervised scenario where no anomaly labels are available. In this case, the pre-training stage remains unchanged. After obtaining the pre-trained low- and high-pass node representations, we directly concatenate them and feed the representations into Isolation Forest (IF) (Liu et al., 2008), a widely used ensemble method for unsupervised anomaly detection, to derive anomaly scores for each node. This modification enables our framework to operate in a label-free manner, aligning with common practice in unsupervised GAD.
We follow the evaluation pipeline of the unsupervised GAD benchmark BOND (Liu et al., 2022), and select the three datasets overlapping with ours: Reddit, Weibo, and DGraph. We include comparisons with two recent state-of-the-art unsupervised baselines, GAD-EBM (Roy et al., 2023) and DiffGAD (Li et al., 2025). Table 5 reports AUROC results, where baseline numbers are taken from the respective papers.
Table 5: Unsupervised GAD results (AUROC $\%$).
Method Reddit Weibo DGraph
GAD-EBM (Roy et al., 2023) 58.5±1.6 93.2±1.8 60.3±2.5
DiffGAD (Li et al., 2025) 56.3±0.1 93.4±0.3 52.4±0.0
IF (raw feats) (Liu et al., 2008) 45.2±1.7 53.5±2.8 60.9±0.7
IF + Our Pre-training 59.9±2.4 93.1±1.8 63.3±0.6
Ours (Semi-supervised) 66.8±3.9 98.8±0.3 72.4±1.3
From the results, we observe that: (i) pretraining substantially improves the performance of IF, which alone struggles to capture structural anomalies; (ii) our method surpasses the state-of-the-art DiffGAD on Reddit and DGraph in the unsupervised setting; (iii) nevertheless, the semi-supervised version of our full framework still achieves the best performance, indicating that anomaly-aware pretraining brings general benefits while labels further boost detection accuracy.
Overall, these findings demonstrate that the proposed pre-training not only benefits semi-supervised settings but also provides clear gains in purely unsupervised anomaly detection, further validating its general effectiveness.
I.2 Visualization of Learned Representations
To provide more intuitive insights into the learned representations, we apply t-SNE to visualize the distributions of $\boldsymbol{Z}_{L}$, $\boldsymbol{Z}_{H}$, and their fused representation $\boldsymbol{Z}$. As shown in Figure 6, the results consistently reveal clear patterns: (i) the low-pass representation $\boldsymbol{Z}_{L}$ and the high-pass representation $\boldsymbol{Z}_{H}$ typically occupy distinct and largely non-overlapping regions in the embedding space, indicating that they capture complementary signals; (ii) the fused representation $\boldsymbol{Z}$ exhibits partial overlaps with both $\boldsymbol{Z}_{L}$ and $\boldsymbol{Z}_{H}$, suggesting that it effectively integrates information from both frequency domains. These visualizations thus provide direct evidence that the dual-filter encoding indeed extracts complementary information, and the adaptive fusion mechanism successfully combines them into more comprehensive node representations.
(a) Reddit
(b) Weibo
(c) Amazon
(d) YelpChi
(e) T-Finance
(f) Elliptic
(g) Tolokers
(h) Questions
Figure 6: Visualization of the learned representations $\boldsymbol{Z}_{L}$, $\boldsymbol{Z}_{H}$, and $\boldsymbol{Z}$.
I.3 Efficiency Comparison.
@@ I.2 Visualization of Learned Representations
I.4 Hyperparameter Analysis
In addition to the adaptive fusion mechanism, our APF further introduces two hyperparameters, $p_{a}$ and $p_{n}$, which represent the expected preference for low-pass representations in abnormal and normal nodes, respectively. To assess their impact on our performance, we vary these hyperparameters from $0.0$ to $1.0$ in increments of $0.1$. The results are presented in Figure 7, 8 and 9. It is observed that the right half of the heatmap, corresponding to relatively larger $p_{n}$ values, generally outperforms the left half. This aligns with the understanding that normal nodes benefit more from low-pass representations for generic knowledge, due to their strong structural consistency with neighbors. Additionally, the optimal combination of $(p_{a},p_{n})$ always appears in the lower-right half of the heatmap, where $p_{a}\leq p_{n}$. This indicates that abnormal nodes are assigned lower $p_{a}$ values, thus placing greater emphasis on anomaly-indicative components, which better capture their deviation from the local context. These observations are consistent with the intuition behind our adaptive fusion design: normal and abnormal nodes require different emphases of knowledge to maximize discriminability. Overall, $p_{a}$ and $p_{n}$ provide a simple yet effective way that consistently guides APF toward strong and stable anomaly detection performance with minimal tuning effort.
(a) Weibo
(b) Amazon
(c) YelpChi
(d) T-Finance
(a) Weibo
(b) Amazon
(c) YelpChi
(d) T-Finance
(a) Weibo
(b) Amazon
(c) YelpChi
(d) T-Finance
I.5 Under Varying Supervision
This section evaluates the performance of APF under varying levels of supervision by modifying the number of labeled abnormal nodes. Following Tang et al. (2023), the number of labeled normal nodes is set to four times the number of labeled abnormal nodes. We present the results in terms of AUPRC, AUROC, and Rec@K in Figures 10, 11, and 12, respectively. As expected, performance generally improves across all methods as the number of labeled nodes increases. Notably, APF delivers consistent improvements over baseline pre-training methods and surpasses the state-of-the-art GAD models, even with only 5 labeled abnormal nodes. This highlights the effectiveness of our approach in addressing GAD with limited supervision.
@@ I.5 Under Varying Supervision
I.6 Additional Results for Model Comparison.
@@
Table 7: Comparison of AUROC for each model. ”-” denotes ”out of memory”. The best and runner-up models are bolded and underlined.
@@
Table 8: Comparison of Rec@K for each model. ”-” denotes ”out of memory”. The best and runner-up models are bolded and underlined.
@@
I.7 Additional Figures for Homophily Disparity.
We provide additional figures in terms of AUROC and Rec@K for homophily disparity in Figure 13 and Figure 14, respectively.
@@
I.8 Performance Analysis by Supervision Paradigm
To clarify performance differences and better position our method within the landscape of GAD approaches, we provide an additional analysis categorizing baseline methods by the type of anomaly information used. While our main text categorizes models by architecture (e.g., Standard GNNs vs. Specialized GAD models), here we classify them based on their reliance on supervision, particularly under the label-scarce setting (100 labeled nodes) used in our experiments:
  • Supervised Models: These models are trained directly using the available labeled anomalies. This category includes standard GNNs (e.g., GCN, GAT, ACM) and supervised GAD-specific models (e.g., GAS, PCGNN, AMNet, BWGNN, GHRN).
  • Semi-supervised Models: These are specialized subclasses of supervised methods designed to effectively leverage limited anomaly labels. This category includes ConsisGAD, SpaceGNN, and XGBGraph.
  • Self-supervised Models: These models adopt a two-stage paradigm: first pre-training on unlabeled data to learn general representations, followed by fine-tuning with labeled anomalies. This category includes general graph pre-training methods (e.g., DGI, DCI, GraphMAE, SSGE, PolyGCL, BWDGI) and our proposed APF.
Table 9: Comparison of top-performing models across different supervision paradigms. Results are averaged across all 10 datasets. Our proposed APF (Self-supervised) demonstrates superior performance compared to the best models in Supervised and Semi-supervised categories.
Category Model Avg. AUPRC Avg. AUROC Avg. Rec@K
Supervised PCGNN 32.9 73.8 34.6
BWGNN 35.4 76.0 37.4
GHRN 35.4 76.1 37.7
Semi-supervised ConsisGAD 38.6 78.0 39.6
SpaceGNN 41.8 79.0 42.2
XGBGraph 42.9 77.8 41.3
Self-supervised SSGE 37.1 77.6 38.2
PolyGCL 37.1 78.7 38.6
BWDGI 39.2 78.2 39.2
APF (Ours) 49.6 82.8 48.4
Table 9 summarizes the performance of the top-3 models from each category alongside APF. This categorization yields two critical observations regarding label scarcity in GAD: